# Compression for Better: A General and Stable Lossless Compression Framework

### v1

@@ 3 Lossless Theoretical Framework
[…]
@@
where $f(\cdot)$ represents the loss of the model on a dataset, $\mathbf{E}$ stands for expectation, $m$ is the size of the dataset, $\ell(\cdot)$ is the loss function for a sample, and $(x_{i},y_{i})$ denotes a sample in the dataset along with its corresponding label, $L(\cdot)$ represents the loss function, such as the cross-entropy function; $h_{i}$, with $i\in[1,...,n]$, represents the ($n-i+1$)th layer in the neural network; $W=(w_{n}^{T},w_{n-1}^{T},\cdots,w_{1}^{T})^{T}$, where $w_{i}$ is the parameter in $h_{i}(\cdot)$; and for the reason of a unified format, $h_{n+1}$ denotes the sample $x$. This form ensures that LLC is independent of the specific network architecture, making it applicable to different models.
@@
¯$\bar{\ell}(w,x_{i},y_{i})=L(h_{1}(h_{2}(\cdots h_{n}(h_{n+1}+\epsilon_{n},w_{n}+\delta_{n})$ (2)
@@
¯$\bar{\ell}(w,x_{i},y_{i})-\ell(w,x_{i},y_{i})=\sum_{i=1}^{n}\frac{\partial\ell}{\partial h_{i+1}}\cdot\epsilon_{i}+\frac{\partial\ell}{\partial w_{i}}\cdot\delta_{i}+$ (3)
$\frac{1}{2}(\epsilon_{i},\delta_{i})^{T}\mathbb{H}(\epsilon_{i},\delta_{i})+O(||(\epsilon_{i},\delta_{i})||^{n})$
where $\mathbb{H}$ represents the Hessian matrix and $O(||(\epsilon_{i},\delta_{i})||^{n})$ represents the high-order term, $\cdot$ is inner product and * is the scalar product. For the loss on whole dataset, we can gain
@@
$+\frac{1}{2}(\epsilon_{i},\delta_{i})^{T}\mathbb{H}(\epsilon_{i},\delta_{i})+O(||(\epsilon_{i},\delta_{i})||^{n})$
@@
When the variations $\epsilon,\delta$ are small enough, the actual change in the loss function can be accurately described by the total differential $df$. The lemma above outlines the theoretical range for noise. Thus, it is essential to identify this "sufficiently small" threshold within the practical model.
@@
the left side represents the loss due to actual noise disturbances, while the right side represents the theoretical loss induced by noise. The $k$ controls the compression level, such as 4/8-bit or rank. Eq. 5 and Eq. 6 calculate the noise bounds for weights and activations, respectively. LLC measures the change in loss through perturbation of the first- and second-order terms. As shown in Fig. 1, the perturbation noise boundaries under LLC and their dominant influencing orders are illustrated. First, for activations, when the noise range is below $10^{-3}$, the first-order term is the dominant factor, and the gradient can be treated as the primary optimization target, with the influence of higher-order terms negligible. This is because higher-order errors decay exponentially compared to lower-order ones. When the noise range is between $[10^{-3},8\times 10^{-2}]$, both the first- and second-order terms significantly affect the loss, thus the second-order Hessian information should be included in the optimization target. For noise levels above $8\times 10^{-2}$, the negative impact increases, and higher-order terms need to be considered.
For weights, ideally, the first-order term in a trained-well model should be zero. However, in practice, gradients are rarely zero, so they must be included in the optimization. When the noise is less than $8\times 10^{-3}$, the first-order term should be the main optimization target. When the noise range is between $[8\times 10^{-3},2\times 10^{-1}]$, the second-order term’s influence should be considered. When the noise exceeds $2\times 10^{-1}$, the influence of weight noise on the loss becomes significant, and higher-order terms should not be omitted. This range defines the noise boundaries and the dominant terms in the compression process.
@@ 4 LLC Quantization and Decomposition
Quantization. Lossless mixed-precision quantization addresses two key challenges: first, how to achieve stable lossless compression under mixed-precision quantization; and second, how to efficiently select the optimal quantization bit-width for each layer, which is an NP-hard problem. For the first challenge, LLC quantization is applied for first-order analysis, ensuring lossless quantization within the first-order bounds. The second challenge is reformulated as a group knapsack problem, which is solved efficiently using dynamic programming. In the LLC framework, the loss function is treated as the "value $P$", each layer $i$ is considered a "group" with one bit-width $j$ choice per group, and the model size is treated as the "knapsack capacity $W$". This transforms the original problem into a low-computation group knapsack problem, where the goal is to select the optimal bit-width for each layer to minimize loss while keeping the quantized model size within the specified capacity $C$.
@@
where $n$ is the number of model layers. The problem scale of the grouped backpack is very small, usually less than $n*k$, and has a significant efficiency advantage. The overall process of our proposed method is shown in Algorithm 1.
@@
1:Input: Neural network $M$ with $n$ layers, quantization levels $[q_{1},q_{2},...,q_{k}]$, maximum error $error_{max}$, calibration dataset $D$
@@
7:         Compute $||\epsilon_{i}||$ and $scale_{input}$ for $Layer_{i}$
8:         Calculate $slope$ = $\frac{||f(M)-f_{input}(M;scale_{input},i)||}{scale_{input}}$
9:         Compute $fluc$ as
10:$||f(M)-f_{weight}(M;scale_{weight},i)||$
@@
18:              Update $P[i][j]$ with $slope\times\frac{||\epsilon_{i}||}{\sqrt{size(e_{i})}}$
@@
20:              Calculate $||\delta_{i}||$ and $scale_{weight}$
21:              Update $P[i][j]$ with $slope\times\frac{||\epsilon_{i}||}{\sqrt{size(\epsilon_{i})}}+\frac{fluc}{scale_{weight}}\times\frac{||\delta_{i}||}{\sqrt{size(\delta_{i})}}$
@@
In the algorithm, $\epsilon$ and $\delta$ are the quantization noise errors of activation and weight. Positive and negative are the choices of different quantization directions. We set the quantization level of Algorithm 1 to $k=4$ categories, namely 2/4/8/16 bit. The total time complexity is $O(n*k*feature)$.
@@
1:Neural network $M$ with $n$ layers, loss threshold $\epsilon$, maximum rank $rank_{max}$
@@
7:     Initialize loss list $A$ for recording candidate factorizations.
@@
9:         Initialize $L_{c}$, $R_{c}$triangleright Temporary matrices for rank-$c$ factorization
@@
13:                  {Record $L_{c}$, $R_{c}$, and computed Loss} in list $A$
14:                  if $\{\|w_{ij}-l_{ij}r_{ij}\|_{F}\}_{i,j}$rightarrow 0)  then
@@
19:         Update $L_{c}$, $R_{c}$triangleright Update
@@
21:     Select $L_{a}$, $R_{a}$ from $A$ that minimizes Loss
22:     $W_{a}=L_{a}\cdot R_{a}$triangleright Final factorized matrix for layer
@@
$\operatorname*{min}_{\delta^{k}\in\Delta}\bar{f}(w)-f(w)$ $=\frac{1}{m}\sum_{i=1}^{n}\sum_{(x_{j},y_{j})\in\mathbb{D}}\frac{\partial\ell}{\partial w_{i}}\cdot\delta^{k}_{i}$ (10)

### v2

@@ 3 Lossless Theoretical Framework
$\begin{aligned} \min_{W}f(W)=\mathbf{E}_{Sample}\ell(W,Sample)=\frac{1}{m}\sum_{(x_{i},y_{i})\in\mathbb{D}}\ell(W,x_{i},y_{i})\end{aligned}$ (1)
@@
where $f(\cdot)$ represents the loss of the model on a dataset, $\mathbf{E}$ stands for expectation, $m$ is the size of the dataset, $\ell(\cdot)$ is the loss function for a sample, and $(x_{i},y_{i})$ denotes a sample in the dataset along with its corresponding label, $L(\cdot)$ represents the loss function, such as the cross-entropy function; $h_{i}$, with $i\in[1,...,n]$, represents the (ni+1n-i+1)th layer in the neural network; $W=(w_{n}^{T},w_{n-1}^{T},\cdots,w_{1}^{T})^{T}$, where $w_{i}$ is the parameter in $h_{i}(\cdot)$; and for the reason of a unified format, $h_{n+1}$ denotes the sample $x$. This form ensures that LLC is independent of the specific network architecture, making it applicable to different models.
@@
$\bar{\ell}(w,x_{i},y_{i})=L(h_{1}(h_{2}(\cdots h_{n}(h_{n+1}+\epsilon_{n},w_{n}+\delta_{n})$ (2)
@@
$\bar{\ell}(w,x_{i},y_{i})-\ell(w,x_{i},y_{i})=\sum_{i=1}^{n}\frac{\partial\ell}{\partial h_{i+1}}\cdot\epsilon_{i}+\frac{\partial\ell}{\partial w_{i}}\cdot\delta_{i}+$ (3)
[…]
@@
[…]
@@
When the variations $\epsilon,\delta$ are small enough, the actual change in the loss function can be accurately described by the total differential dfdf. The lemma above outlines the theoretical range for noise. Thus, it is essential to identify this "sufficiently small" threshold within the practical model.
@@
the left side represents the loss due to actual noise disturbances, while the right side represents the theoretical loss induced by noise. The $k$ controls the compression level, such as 4/8-bit or rank. Eq. 5 and Eq. 6 calculate the noise bounds for weights and activations, respectively. LLC measures the change in loss through perturbation of the first- and second-order terms. As shown in Fig. 1, the perturbation noise boundaries under LLC and their dominant influencing orders are illustrated. First, for activations, when the noise range is below $10^{-3}$, the first-order term is the dominant factor, and the gradient can be treated as the primary optimization target, with the influence of higher-order terms negligible. This is because higher-order errors decay exponentially compared to lower-order ones. When the noise range is between $10^{-3},8\times 10^{-2}]$, both the first- and second-order terms significantly affect the loss, thus the second-order Hessian information should be included in the optimization target. For noise levels above $8\times 10^{-2}$, the negative impact increases, and higher-order terms need to be considered.
For weights, ideally, the first-order term in a trained-well model should be zero. However, in practice, gradients are rarely zero, so they must be included in the optimization. When the noise is less than $8\times 10^{-3}$, the first-order term should be the main optimization target. When the noise range is between $8\times 10^{-3},2\times 10^{-1}]$, the second-order term’s influence should be considered. When the noise exceeds $2\times 10^{-1}$, the influence of weight noise on the loss becomes significant, and higher-order terms should not be omitted. This range defines the noise boundaries and the dominant terms in the compression process.
@@ 4 LLC Quantization and Decomposition
Quantization. Lossless mixed-precision quantization addresses two key challenges: first, how to achieve stable lossless compression under mixed-precision quantization; and second, how to efficiently select the optimal quantization bit-width for each layer, which is an NP-hard problem. For the first challenge, LLC quantization is applied for first-order analysis, ensuring lossless quantization within the first-order bounds. The second challenge is reformulated as a group knapsack problem, which is solved efficiently using dynamic programming. In the LLC framework, the loss function is treated as the "value PP", each layer $i$ is considered a "group" with one bit-width $j$ choice per group, and the model size is treated as the "knapsack capacity WW". This transforms the original problem into a low-computation group knapsack problem, where the goal is to select the optimal bit-width for each layer to minimize loss while keeping the quantized model size within the specified capacity $C$.
@@
where $n$ is the number of model layers. The problem scale of the grouped backpack is very small, usually less than nkn*k, and has a significant efficiency advantage. The overall process of our proposed method is shown in Algorithm 1.
@@
1:Input: Neural network $M$ with $n$ layers, quantization levels $q_{1},q_{2},...,q_{k}]$, maximum error $error_{max}$, calibration dataset $D$
@@
7:   Compute $\epsilon_{i}||$ and $scale_{input}$ for $Layer_{i}$
[…]
9:   Compute flucfluc as
[…]
@@
[…]
@@
20:     Calculate $\delta_{i}||$ and $scale_{weight}$
[…]
@@
In the algorithm, $\epsilon$ and $\delta$ are the quantization noise errors of activation and weight. Positive and negative are the choices of different quantization directions. We set the quantization level of Algorithm 1 to $k=4$ categories, namely 2/4/8/16 bit. The total time complexity is O(nkfeature)O(n*k*feature).
@@
1:Neural network $M$ with $n$ layers, loss threshold $\epsilon$, maximum rank $rank_{max}$ .
@@
7:  Initialize loss list AA for recording candidate factorizations.
@@
9:   Initialize $L_{c}$, $R_{c}$ $\triangleright$ Temporary matrices for rank-$c$ factorization
@@
13:      {Record $L_{c}$, $R_{c}$, and computed Loss} in list AA
14:      if $\{\|w_{ij}-l_{ij}r_{ij}\|_{F}\}_{i,j}$ $\rightarrow$ 0) then
@@
19:   Update $L_{c}$, $R_{c}$ $\triangleright$ Update
@@
21:  Select $L_{a}$, $R_{a}$ from AA that minimizes Loss
22:  $W_{a}=L_{a}\cdot R_{a}$ $\triangleright$ Final factorized matrix for layer
@@
$\operatorname*{min}_{\delta^{k}\in\Delta}\bar{f}(w)-f(w)$ $=\frac{1}{m}\sum_{i=1}^{n}\sum_{(x_{j},y_{j})\in\mathbb{D}}\frac{\partial\ell}{\partial w_{i}}\cdot\delta^{k}_{i}\ \$(10)